Find an equatin of the line that contains the given pair of points.
(22, 25), (5,8)

Answer by lwsshak3(11628)

Find an equatin of the line that contains the given pair of points.
(22, 25), (5,8)
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
For given equation:
slope,m=∆y/∆x=rise/run=(25-8)/(22-5)=17/17=1
y=(1)x+b
solve for b using any one of the points (5,8)
8=5+b
b=3
equation:
y=x+3